f={(1,2),(2,3),(3,3)},g={(1,0)(2,3),(3,2),(4,8)}
Find the range of f+g=

Answer by greenestamps(13334)



f+g is only defined for values of x in the domains of both f and g: {1,2,3}. f+g is not defined for x=4, because 4 is in the domain of only g, not f.

We aren't going to run into any problems with function values being undefined for the function f+g, so there will be three values in the range of f+g.

(f+g)(1) = 2+0 = 2
(f+g)(2) = 3+3 = 6
(f+g)(3) = 3+2 = 5

ANSWER: The range of f+g is {2,5,6}